OpenStudy (anonymous):

What is the slope of a line with the points (41, 33) and (62, 27)? A.) zero B.) positive C.) undefined D.) negative

OpenStudy (anonymous):

\[\frac{\Delta y}{\Delta x}=\frac{33-27}{41-62}=m=slope\]

OpenStudy (ybarrap):